OpenStudy (superhelp101):

When 3.5 mol Al react with 12.5 mol HCl, what is the limiting reactant and how many moles of H2 can be formed? 2 Al + 6 HCl 2 AlCl3 + 3 H2 HCl is the limiting reactant, 6.3 mol H2 can be formed HCl is the limiting reactant, 3.2 mol H2 can be formed Al is the limiting reactant, 3.5 mol H2 can be formed Al is the limiting reactant, 5.3 mol H2 can be formed

OpenStudy (superhelp101):

@Festinger

OpenStudy (festinger):

Have you found the amount of mols of each reactant?

OpenStudy (festinger):

Oh right it's given. Compare the ratios.

OpenStudy (festinger):

\[amount\:Al\:needed\:if\:HCl\:is\:added:in\:excess=(\frac{mol\:of\:Al}{mol\:of\:HCl})*mols\:of\:HCl\] \[amount\:HCl\:needed\:if\:Al\:is\:added:in\:excess=(\frac{mol\:of\:HCl}{mol\:of\:Al})*mols\:of\:Al\]

OpenStudy (festinger):

So if I had Al in excess, the calculation would look like: \[\frac{6}{2}*3.5=10.5mols\:of\:HCl\] In other words, 10.5mols of HCl reacts with 3.5mols of Al.

OpenStudy (festinger):

How many mols of Al are needed to react with 12.5mols of HCl?

OpenStudy (superhelp101):

ok hold i going to work it out

OpenStudy (superhelp101):

do i have to use the first formula you provide me to solve the question : How many mols of Al are needed to react with 12.5mols of HCl?

OpenStudy (festinger):

Yeap

OpenStudy (superhelp101):

\[\frac{ 3 }{ 12.5 }time12.5\]

OpenStudy (superhelp101):

i am stuck

OpenStudy (festinger):

nope, the term in brackets is the ratio from the balanced equation

OpenStudy (superhelp101):

oh so i used the wrong formula

OpenStudy (superhelp101):

\[\frac{ 6 }{ 2 }\times variable =12.5\]

OpenStudy (festinger):

\[\frac{2}{6}\times 12.5=\frac{25}{6}\]

OpenStudy (festinger):

What need need to do is to compare ratios... 6 mols of HCl reacts with 2 mols of Al, so how much Al will be reacted with 12.5mols of HCl?

OpenStudy (superhelp101):

i set up a proportion so i got 4.17

OpenStudy (festinger):

Yes

OpenStudy (superhelp101):

:) is that it?

OpenStudy (festinger):

So do you have enough Al to react with all 12.5 mols of HCl?

OpenStudy (superhelp101):

yes

OpenStudy (festinger):

You only have 3.5 mols of Al, but 12.5 mols of HCl needs 4.16 mols of Al to react fully.

OpenStudy (superhelp101):

what does it mean when it says limited reactant?

OpenStudy (festinger):

It means the reactant that does not have enough (in mols) to react fully Let's say an reaction is \[A+B\rightarrow C\] And I have 27 mols of A, and 5 mols of B. The ratio of A to B is 1:1 Therefore for the reaction to complete fully, 27 mols of A must react with 27 mols of B, while for B we need 5 mols of B to react with 5 mols of A. here, we see that B does not have enough to react with all of A, and we call B the limiting reagent.

OpenStudy (festinger):

In my example, when the reaction completes, I will have 22 mols of A (left over), 0 mols of B and 5 mols of C.

OpenStudy (superhelp101):

so it Al the limiting reagent?

OpenStudy (festinger):

Yes

OpenStudy (festinger):

Now you have to find how much H2 gas is produced from 3.5mols of Al!

OpenStudy (superhelp101):

what formula should i use?

OpenStudy (festinger):

It's similar to the ones we've used before, you just need to modify it abit. Give it a try.

OpenStudy (superhelp101):

5.3

OpenStudy (festinger):

Yes, 5.25 mols of H2 gas are produced. You got it!

OpenStudy (superhelp101):

yay thank you very much for all you help!